# Structured Pruning for Diverse Best-of- Reasoning Optimization

### v1

@@ 1 Introduction
At the same time, recent research on model pruning has shown that comparable performance can be achieved by selectively disabling certain heads in the transformer architectureĀ Ma etĀ al. (2023); Wang etĀ al. (2019). Removing an attention head does not necessarily weaken the overall ability of the model because a head may introduce noise into the decision-making process, leading to incorrect answers. In fact, our empirical observations suggest that pruning certain heads can improve the reasoning methodsā performance on specific categories of questions; SectionĀ 3.2 discusses this phenomenon in more detail. This finding motivates our approach: we aim to construct a pool of pruned models from one base model with diverse reasoning capabilities without incurring additional computational or memory overhead.
Contributions. We begin by analyzing and providing new insights into the surprising phenomenon that pruning attention heads can improve the performance of the language model on the MATH benchmark. Motivated by these insights, we introduce the Structured PRunIng for diverSe reasoNing opTimization (SPRINT) framework. SPRINT first learns a set of embeddings for each layer-attention head pair using contrastive learning, guided by a novel Diversity-Promoted Contrastive Loss. This loss function encourages the question embedding to move closer to the embeddings of heads whose pruning leads to correct answers for that question. Finally, we leverage SPRINT to enhance best-of- sampling by pruning the top nearest layer-model configurations, determined by the distance between the learned embeddings and the question embedding.
@@ 3 Does Pruning an Attention Head Improve LMās Performance?
Attention head pruning. In our study, we define head pruning as zeroing out the output vector of a specific attention head before the concatenation step, as illustrated in FigureĀ 1. In modern decoder-only architecturesĀ , this is equivalent to pruning the input of layer.
@@ 3.1 Experimental Procedure
Layer and Head Selection: We select layers from the model, yielding a total of attention heads. For each head, we create a variant of the base model with that specific head pruned, resulting in model variants in addition to the base model. We denote this model pool as .
@@
In this experiment, we study two models specialized for mathematical reasoning, Qwen2.5-Math-1.5B-Instruct and Qwen2.5-Math-7B-Instruct, as well as two general-purpose models, Qwen2.5-7B-Instruct, and Meta-Llama-3-8B-Instruct. Our experiments involve evaluating model performance with and without head pruning on 500 samples from the MATH500 datasetĀ Hendrycks etĀ al. (2021).
@@ 3.2 Empirical Observations
Our empirical results, shown in TablesĀ 2, 3, 4, and 1, reveal several important and sometimes surprising findings about attention head pruning. Contrary to the common belief that all attention heads are essential for optimal model performance, we find that pruning individual heads often does not reduce accuracy, and in many cases, actually increases it. The impact of pruning depends strongly on both the specific head and its layer: while some heads are crucial (with pruning causing sharp drops in accuracy), many others are unnecessary, and removing them can leave performance unchanged or even lead to notable improvements. For instance, in the āCounting & Probabilityā category, pruning head 0 in layer 15 of Qwen2.5-Math-1.5B-Instruct increased accuracy by 16% over the unpruned baseline.
We further illustrate this trend in FigureĀ 2, which displays violin plots of additive performance gains (the difference between the best pruned head accuracy and the baseline accuracy) across all subcategories for each model. These plots show that head pruning consistently produces positive gains for all evaluated models. The specialized mathematical models (Qwen2.5-Math-1.5B-Instruct and Qwen2.5-Math-7B-Instruct) exhibit higher median gains and more frequent large improvements, particularly in domains such as āCounting & Probabilityā and āGeometryā. In contrast, the general-purpose models (Qwen2.5-7B-Instruct and Meta-Llama-3-8B-Instruct) show more modest but still positive and consistent benefits from pruning.
@@ 4 Methodology
To further quantify the correspondence between different pruned-head models, we introduce a similarity measure between any two heads and as follows:
@@
Our joint optimization loss consists of two components. The first term aims to align the question embedding with the embeddings of those pruned-head models that yield the correct answer. The loss employs a softmax formulation for each question, over the negative squared Euclidean distances between and the head embeddings. This loss encourages the model to produce a lower distance, and thus a higher similarity, for those heads in the positive set , which results in a correct answer.
@@
Overall, our unified methodology demonstrates that pruning the attention heads, guided by a contrastive alignment strategy, can improve model performance. Thus, the SPRINT framework identifies which heads benefit from pruning and supports a dynamic simulation approach to generate optimal predictions.
@@ 5 Numerical Experiments
Base Models. As a proof of concept, we select Qwen2.5-Math-1.5B-Instruct as our generation model. This choice is motivated by its strong empirical gains observed in FigureĀ 2. All experiments are performed using the original, unquantized version of this model. We consider five layers {5, 10, 15, 20, 25} for pruning.
Performance metrics. To assess performance, we use the Pass@ metric. Specifically, each method is used to sample candidate answers, and the accuracy is computed based on the best answer as determined by an oracle.
@@
The projectās repository is available at https://github.com/HieuNT91/attention_pruning.
@@ 5.1 Quantitative Results
FigureĀ 4 summarizes our primary findings. Our method consistently outperforms the baselines for nearly all values of across the MATH and GSM8K datasets. Notably, the greatest improvements are observed at lower values , while the performance gains become more gradual as increases.
@@ 6 Conclusion
In this work, we demonstrated that pruning attention heads, a technique traditionally viewed as harmful, can enhance the reasoning capabilities of transformer-based language models. Our study revealed that the selective removal of specific heads mitigates the propagation of redundant or noisy signals and enhances overall model performance. We developed a contrastive learning framework to harness this phenomenon that dynamically selects the optimal head and layer to prune at inference time. Experimental results confirm that our approach significantly outperforms traditional Best-of- methods and random head selection strategies.
@@
Due to the space limitations of a short paper, our work does not experiment with a more diverse range of model architectures and sizes. Our work primarily focuses on the mathematical domain, but it can be extended to other scientific domains that require LMsā reasoning.
@@
Datasets. The MATH dataset contains 12,500 math problems, with 12,000 problems designated for training and 500 for testing. Similarly, the GSM8K dataset comprises 8,792 math problems, with 7,473 used for training and 1,319 for testing. In our experiments, we subsample 1,500 training examples from the respective training sets. The head embeddings are then trained using the subsampled training set. Finally, we evaluate performance on the test splits of the aforementioned datasets.
@@
FigureĀ 5 shows that pruning an attention head in Qwen-2.5-1.5B Math Instruct model mostly leads to improvement in accuracy@1 on the MATH500 dataset.

### v2

@@ 1 Introduction
At the same time, recent research on model pruning has shown that comparable performance can be achieved by selectively disabling certain heads in the transformer architectureĀ Ma etĀ al. (2023); Wang etĀ al. (2019). Removing an attention head does not necessarily weaken the overall ability of the model because a head may introduce noise into the decision-making process, leading to incorrect answers. In fact, our empirical observations suggest that pruning certain heads can improve the reasoning methodsā performance on specific categories of questions; SectionĀ 3.2 discusses this phenomenon in more detail. This finding motivates our approach: starting from a single base model, we create a set of new models by pruning a different attention head in each, one at a time. This allows us to build a diverse pool of pruned models without significant additional computational or memory overhead.
Contributions. We begin by analyzing and providing new insights into the surprising phenomenon that pruning attention heads can improve the performance of the language model on the MATH benchmark. Motivated by these insights, we introduce the Structured PRunIng for diverSe reasoNing opTimization (SPRINT) framework. SPRINT first learns a set of embeddings for each attention head-layer pair using contrastive learning, guided by a novel Diversity-Promoted Contrastive Loss. This loss function encourages the question embedding to move closer to the embeddings of heads whose pruning leads to correct answers for that question. Finally, we leverage SPRINT to enhance best-of- sampling by pruning the top nearest head-layer configurations, determined by the distance between the learned embeddings and the question embedding.
@@ 3 Does Pruning an Attention Head Improve LMās Performance?
Attention head pruning. In our study, we define head pruning as zeroing out the output vector of a specific attention head before the concatenation step, as illustrated in FigureĀ 1. In modern decoder-only architecturesĀ AI@Meta (2024), this is equivalent to pruning the input of layer.
@@ 3.1 Experimental Procedure
Layer and Head Selection: We select layers from the model, specifically, the first layer, layers 5 and 15, and the last layer, for a total of attention heads. For each head, we create a variant of the base model with that specific head pruned, resulting in model variants in addition to the base model.
@@
In this experiment111The projectās repository is available at https://github.com/HieuNT91/attention_pruning., we study two models specialized for mathematical reasoning, Qwen2.5-Math-1.5B-Instruct and Qwen2.5-Math-7B-Instruct, as well as two general-purpose models, Qwen2.5-7B-Instruct, and Meta-Llama-3-8B-Instruct. Our experiments involve evaluating model performance with and without head pruning on 500 samples from the MATH500 datasetĀ Hendrycks etĀ al. (2021).
@@ 3.2 Empirical Observations
Our empirical results, shown in TablesĀ 1, 2, 3, 4, reveal several surprising findings about attention head pruning. Contrary to the common belief that all attention heads are essential for optimal model performance, we find that pruning individual heads often does not reduce accuracy, and in many cases, increases it. The impact of pruning depends strongly on both the specific head and its layer: while some heads are irreplaceable, for example, removing head 3 in layer 0 (see tablesĀ 2, 3, and 4) causes performance to drop to near zero. This pattern mirrors findings fromĀ Yu etĀ al. (2024), which show that certain āsuperweightsā in the early layers play a disproportionate role in overall LLM performance. Conversely, pruning less critical heads can preserve or even enhance performance. For instance, in the āCounting & Probabilityā category, pruning head 0 in layer 15 of Qwen2.5-Math-1.5B-Instruct increased accuracy by 16% over the unpruned baseline.
We further illustrate this trend in FigureĀ 2, which displays violin plots of additive performance gains (the difference between the best pruned head accuracy and the baseline accuracy) across all subcategories for each model. These plots show that head pruning consistently produces positive gains for all evaluated models. The specialized mathematical models exhibit higher median gains and more frequent large improvements, particularly in domains such as āCounting & Probabilityā and āGeometryā. In contrast, the general-purpose models show more modest but still positive and consistent benefits from pruning.
@@ 4 Methodology
To quantify the correspondence between different pruned-head models, we introduce a similarity measure between any two heads and as follows:
@@
Our joint optimization loss consists of two components. The first term aims to align the question embedding with the embeddings of those pruned-head models that yield the correct answer. The loss employs a softmax formulation for each question, over the negative squared Euclidean distances between and the head embeddings. This loss encourages the model to produce a lower distance, and thus a higher similarity, for those heads in the positive set .
@@
Overall, our unified methodology demonstrates that pruning the attention heads, guided by a contrastive alignment strategy, can improve model performance. Thus, SPRINT identifies which heads benefit from pruning and supports a dynamic simulation approach to generate optimal predictions.
@@ 5 Numerical Experiments
Base Models. As a proof of concept, we select Qwen2.5-Math-1.5B-Instruct as our generation model. This choice is motivated by its strong empirical gains observed in FigureĀ 2. All experiments are performed using the original, unquantized version of this model. Due to resource constraints, we consider five layers {5, 10, 15, 20, 25} for pruning.
Performance metrics. To evaluate performance, we use the Pass@ metric. For each method, we sample candidate answers and compute Pass@ based on whether at least one candidate is correct, as determined by an oracle.
@@ 5.1 Quantitative Results
FigureĀ 4 summarizes our primary findings. Our method consistently outperforms the baselines for nearly all values of across the MATH and GSM8K datasets. Notably, the greatest improvements are observed at lower values , while the performance gains become more gradual as .
@@ 6 Conclusion
In this work, we demonstrated that pruning attention heads, a technique traditionally viewed as harmful, can enhance the reasoning capabilities of transformer-based language models. Our study revealed that the selective removal of specific heads mitigates the propagation of redundant or noisy signals, thereby enhancing overall model performance. We developed a contrastive learning framework to harness this phenomenon that dynamically selects the optimal attention head-layer configurations to prune at inference time. Experimental results confirm that our approach significantly outperforms traditional Best-of- methods and random head selection strategies.
@@
Due to the space limitations of a short paper, our work does not experiment with a more diverse range of model architectures and sizes. Our work primarily focuses on the mathematical domain, but it can be extended to other scientific domains that require large language modelsā reasoning. Furthermore, we did not study an interesting case where pruning a single head leads to a catastrophic performance drop.
@@
Datasets. The MATH dataset contains 12,500 math problems, with 12,000 problems designated for training and 500 for testingĀ Lightman etĀ al. (2023). Similarly, the GSM8K datasetĀ Cobbe etĀ al. (2021) comprises 8,792 math problems, with 7,473 used for training and 1,319 for testing. In our experiments, we subsample 1,500 training examples from the respective training sets. The attention head embeddings are then trained using the subsampled training set. Finally, we evaluate performance on the test splits of the aforementioned datasets.
@@
FigureĀ 5 shows that pruning an attention head in Qwen-2.5-1.5B-Math-Instruct model mostly leads to improvement in accuracy on the MATH500 dataset.